hi how do you solve this:
write an equation for the parabola with the focus (1,3) and vertex (0,3)

Answer by lwsshak3(11628)

hi how do you solve this:
write an equation for the parabola with the focus (1,3) and vertex (0,3)
**
This is a parabola that opens rightwards: (focus to right of vertex)
Its standard form of equation: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex
For given parabola:
Axis of symmetry: y=3
p=1 (distance from focus to vertex on the axis of symmetry)
4p=4
Equation: (y-3)^2=4x